A total of $30,000 was invested in two mutual funds. One fund yields 10.5% annual simple interest; the other fund earns 6%. At the end of the year, the combined interest from the two funds was $2,340. How was invested in each fund. Thank You
Answer by checkley77(12844)

.105X+.06(30,000-X)=2,340
.105X+1,800-.06X=2,340
.045X=2,340-1,800
.045X=540
X=540/.045
X=12,000 AMOUNT INVESTED @ 10.5%
30,000-12,000=18,000 AMOUNT INVESTED @ 6%
PROOF:
.105*12,000+.06*18,000=2,340
1,260+1,080=2,340
2,340=2,340